This problem will require you to think about gene expression (transcription and translation) as well as energy metabolism. 1. Consider a yeast cell which wants to express a 900 bp gene. Assuming this gene has no introns, ignoring any 5’ or 3’ untranslated regions, and assuming a 100 bp poly-A tail, how many nucleotides are required to transcribe this gene? For simplicity’s sake, let’s assume all the nucleotides are ATP. 2. Based on what you know about translation, how many ATPs (or GTPs) would required to translate this protein? 3. Given the total number of ATPs required for transcription and translation, how many glucose molecules would be required to generate enough ATPs? 4. If the total amount of protein in the yeast cell is 10-20 mol, how much glucose is required to synthesize all the cellular protein (assume all proteins are identical to the one you’ve calculated in steps 1-3)?
Answer #1

No. of base pairs in the given gene= 900 bp

All the 900 bp region is the coding region, therefore for its transcription, 900 nucleotides are required, but to add a tail of 100 bp, we need 100 more nucleotides. therefore we require 1000 bases (or ATPs) to synthesize the RNA.

In translation, 2 GTP molecules (say ATP) are required for each codon (3 bases), therefore for 900 bases of gene we require 2 x 300 = 600 ATP molecules

Total ATP required = 1000 + 600 = 1600 ATP molecules

1 mol. of glucose on complete oxidation yields 36 ATP molecules

therefore no. of glucose molecules required to produce 1600 ATP molecules = 1600/36 =45 glucose molecules

For 1 protein molecule we require 45 glucose molecules

for 10 molecules= 45 x 10= 450

For 20 molecules= 45x 20= 900

it means 450- 900 glucose molecules.